Circular No. 6774

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

X1354–644 = GS 1354–64 B. A. Harmon and C. R. Robinson, Marshall Space Flight Center,

NASA, report for the BATSE/Compton Observatory Science Team: “This x-ray transient (cf. IAUC 6772; Kitamoto et al. 1990, Ap.J. 361, 590) is being detected in hard x-rays with emission up to 200 keV — the outburst being first visible to BATSE around Oct. 29 at a level of 60 mCrab (20– 100 keV), where it remained until Nov. 7. More recent data indicate that the source intensity (Nov. 12–18) has increased to 160 mCrab, considerably higher than the reported 2–10-keV flux from RXTE (IAUC 6772). Its location is consistent with Ginga source GS 1354–64, Cen X-2, and MX 1353–64. Evolution from a hard spectrum (power-law index −2.1 to −2.3) to a softer one (index −2.7) occurred between early November and Nov. 18, possibly with a cutoff above 100 keV.”

PKS 2155–304 P. Sreekumar, Goddard Space Flight Center; and W. T. Vestrand, Uni-

versity of New Hampshire, on behalf of the PKS 2155–304 multiwavelength campaign, report: “Quick-look analysis of data taken during Nov. 11–17 by the Energetic Gamma Ray Experiment Telescope aboard the Compton Gamma Ray Observatory indicates that this variable γ-ray blazar is currently displaying an outburst of γ-ray emission; the current flux > 100 MeV is roughly a factor of three greater than the highest flux previously seen in PKS 2155–304. Multiwavelength observations of PKS 2155–304 are urged by southern-hemisphere observers during the outburst. Additional γ-ray observations are scheduled for Nov. 18–24 and Dec. 30–Jan. 13.”

SUPERNOVA 1997ds IN MCG –01-57-007 F. Patat and H. Boehnhardt, European Southern Observatory; and

X. Delfosse, Geneva Observatory, report that a preliminary reduction of a spectrogram (range 350-890 nm, resolution 1.1 nm FWHM) of SN 1997ds, obtained with the 1.54-m Danish telescope (+ DFOSC) on Nov. 20.01 UT at La Silla, shows Hα and Hβ lines with well-developed P-Cyg profiles and expansion velocities (deduced from the absorption minima, uncorrected for the galaxy recession) of 9600 and 7500 km/s, respectively. The Na I 589.2nm line has a marked P-Cyg profile (expansion velocity 7400 km/s). The emission peaks of H lines indicate a recession velocity of 1700 km/s for the parent galaxy. The blue continuum and these other features are consistent with a type-II classification, ∼ 2 weeks past maximum light.

1997 November 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green